1. What is the volume occupied by 0.248 mol of helium gas at 1.33 atm and 301 K?

2. A cylinder contains 30.5 L of oxygen gas at a pressure of 1.9 atm and a temperature of 288 K. How many moles of gas are in the cylinder?

3.

An experiment shows that a 256 mL gas sample has a mass of 0.432 g at a pressure of 745 mmHg and a temperature of 28 ∘C.

What is the molar mass of the gas?
Answer #1

1) Using ideal gas equation :

V = nRT/P = 0.248×0.082057×301/1.33 L

= 4.6 L

2) n = PV/RT = 1.9×30.5/0.082057×288 moles

= 2.45 moles

3) 760 mmHg = 1atm

745 mmHg = 745/760 atm = 0.98 atm

T = 28°C = 28+273.15 K = 301.15 K

V = 256 ml = 0.256 L

Molar mass = mass×RT/PV

= 0.432×0.082057×301.15/0.98×0.256 g/mol

= 42.55 g/mol

Please give feedback!